Calculate the H3O+ concentration and then pH in a 0.100 M HC2H3O2 solution that is 1.34% ionised?


1
Expert's answer
2021-01-25T04:30:47-0500

HC2H3O2 ≡ CH3COOH - acetic acid


Solution:

HC2H3O2(aq) + H2O(l) ⇌ H3O+(aq) + C2H3O2(aq)

ICE Table:



[H3O+] = [C2H3O2-] = αC (according to the ICE Table).

α = 1.34% = 0.0134

C = 0.100 M

Hence,

[H3O+] = αC = 0.0134 × 0.100 M = 1.34×10-3 M

[H3O+] = 1.34×10-3 M


pH is defined as the negative logarithm of [H3O+]:

pH = -log[H3O+]

pH = -log(1.34×10-3) = 2.87

pH = 2.87


Answer: [H3O+] = 1.34×10-3 M; pH = 2.87